OpenStudy (anonymous):

solve the equation: sin x + cos x= =sqrt2 on the interval 0< x <2pi

OpenStudy (anonymous):

I think the first step is to square both sides, so \[\sin x + \cos x = -\sqrt{2}\] becomes \[\sin ^{2}x + \cos ^{2}x = -2\]

OpenStudy (anonymous):

I have two identities that I think apply: \[\sin ^{2}x + \cos ^{2} =1\] and \[\sin ^{2}x = (1-2cosx)/2\]

OpenStudy (raden):

the equation likes sin x + cos x = sqrt(2) or sin x + cos x = -sqrt(2) ?

OpenStudy (anonymous):

The second one, sorry. Had a typo in my original post.

OpenStudy (cwrw238):

i'm afraid sin^2x + cos^x is not the square of sin x + cos x so that method won't work one way to do this is by the complementary angle method

OpenStudy (cwrw238):

- give me a little time to remind myself of this...

OpenStudy (cwrw238):

rather its called the auxiliary angle method but i can't find my notes on it

OpenStudy (tamtoan):

do this, square both sides of the equation, you have : sin^2(x) + 2sin(x)cos(x) + cos^2(x) = 2 and since sin^2(x) + cos^2(x) = 1 , 2 sin(x)cos(x) = sin(2x) you will have sin(2x) + 1 = 2...so, sin(2x) = 1...can you solve 2x from here ? and then x ?

OpenStudy (tamtoan):

i can't use the math notation, so it's a little bit hard to read :), if you have further question, please ask :)